• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 08:02:00 on November 5, 2005

    Entry number 159179

    keyword=Owl Shift Summary

    Owl Shift for Saturday November 5th, 2005


    Leader:   Bryan Moffit
    Worker/Target:   Richard Wilson

    Left HRS p0   = -3.1234 GeV/c   Septum Set Current = 323.61A
    Right HRS p0  = -3.1126 GeV/c   Septum Set Current = 316.92A
    Sweeper is ON at 111.00A

    Summary

    All production this shift (production runs 6168-6176).
    Jaideep's perl script gives me these numbers (removing 6175):
     Total Number of ok Pairs = 304.4 kPairs
     Average beam current = 56.9 µA
     Total Coulombs = 1.125 C
    

    Significant Events

    • 00:00 Began shift near the start of run 6168, compton run 12687. IHWP OUT.
    • 00:02 Incoming Call: T. Holmstrom, status check
    • 00:45 Made a request to MCC to update the Hall-A General Info MEDM screen to display the correct EPICS variables under "Current Status". All of them currently readback "Down" even though there is beam in Hall-A and Hall-B.
    • 04:18 Lost comm with left arm magnet IOC. Reboot through hareboot10. Comm returns within 4 minutes. Cycled Q2 and Q3, before setting p0.
    • 04:58 Had trouble getting Left Q2 back into negative polarity. Paged Tim... asked for Controlled access... paged Mark. Just before Mark called in, polarity of Q2 changed to negative (sorry Mark). Cancelled the controlled access request, and we're back to 57 µA.
    • 06:27 Lost comm with Septum IOC, in the middle of a profile scan. Scan appears to be fine.. Asked MCC to turn off beam during the IOC reboot (Sweeper resets to 0A during reboot... that's bad when there's beam).
    • 06:35 Incoming Call: T. Holmstrom, status check
    • 06:40 Septum IOC recovered. Septa still had current.. had to change left from 323.80 to 323.61A. MCC returned Sweeper to 111A.
    • 07:13 Incoming Call: T. Holmstrom, coffee check
    • 07:29 Tim arrives with the coffee.
    • 00:??

    • Data Run Summary

      • 23:51 Run 6168 - Production Run (20cm LH2), IHWP OUT, 40.2k pairs.
      • 00:47 Run 6169 - Production Run (20cm LH2), IHWP OUT, 37.9k pairs.
      • 01:43 Run 6170 - Production Run (20cm LH2), IHWP OUT, 37.2k pairs.
      • 02:39 Run 6171 - Production Run (20cm LH2), IHWP OUT, 36.7k pairs.
      • 03:35 Run 6172 - Production Run (20cm LH2), IHWP OUT, 41.0k pairs. Left Q2/Q3 down around 84k, due to IOC reboot.
      • 04:30 Run 6173 - Production Run (20cm LH2), IHWP OUT, 32.6k pairs. Left arm recovering... good at about 60k
      • 05:26 Run 6174 - Production Run (20cm LH2), IHWP OUT, 42.3k pairs.
      • 06:22 Run 6175 - Profile Scan (20cm LH2).
      • 06:42 Run 6176 - Production Run (20cm LH2), IHWP OUT, 36.4k pairs.
      • ??:?? Run 6177 - Production Run (20cm LH2), IHWP OUT, Passed to day shift.

      Compton Data Run Summary

      • 23:53 Run 12687 - IHWP OUT. Beam polarization = 86.8%,
      • 02:09 Run 12688 - IHWP OUT. Analysis failed? ROC5 stopped responding.
      • 04:07 Run 12689 - IHWP OUT. Beam polarization = 87.5%, ROC2 stopped responding.
      • 06:55 Run 12690 - IHWP OUT. Beam polarization = 86.6%, ROC0 stopped reporting status.
      • 07:29 Run 12691 - IHWP OUT. Passed to day shift.